OpenStudy (hiroromon):

How do I find the exact value of tan 22.5 degrees?

OpenStudy (anonymous):

you got a calculator?

OpenStudy (hiroromon):

mhm yes i do

OpenStudy (anonymous):

there ya go generally, sin cos and tan of 30, 60, and 90 degrees are suppose to be memorized but any other degree, just grab a calculator

OpenStudy (hiroromon):

wait so i just input that into the calculator?

OpenStudy (anonymous):

oh, one last thing, make sure your calculator is in degrees and not radians

OpenStudy (anonymous):

and yes you jsut input that into a calculator

OpenStudy (anonymous):

well depending on the calculator

OpenStudy (hiroromon):

it cant be decimals though

OpenStudy (anonymous):

why not?

OpenStudy (hiroromon):

because my teacher wants it to be exact, so like in radicals and stuff

OpenStudy (jdoe0001):

are you doing trig?

OpenStudy (anonymous):

tan(x/2) = ±sin(x)) / (1 + cos(x))

OpenStudy (jdoe0001):

hehe

OpenStudy (hiroromon):

do u know how to do it step by step?

OpenStudy (anonymous):

i hate trig identities

OpenStudy (jdoe0001):

@Hiroromon 22.5*2 = 45, thus "x" in the identity will be 45

OpenStudy (jdoe0001):

and 45degrees is the simplest of the better known angles

OpenStudy (anonymous):

using half angle identity tan(x/2) = ±sin(x)) / (1 + cos(x)) tan(x/2) = tan (22.5 degrees) means x/2= 22.5 degrees means x =45 degrees means tan(22.5 degrees) = ±sin(45 degrees )) / (1 + cos(45 degrees))

OpenStudy (hiroromon):

ok ...